OpenStudy (richie63):

EASY! I will Give best Response & Medal! ^.^ 1. Solve. 4y = 144 (Points : 1) 36 16 –16 –36

OpenStudy (cookiemonster2412):

divide both sides by 4

OpenStudy (cookiemonster2412):

144/4=36

OpenStudy (cookiemonster2412):

y=36

OpenStudy (richie63):

^.^ Thanks! :)

OpenStudy (richie63):

Solve. –40 = 5x (Points : 1) –5 –8 –35 –200

OpenStudy (anonymous):

144/4=36 y=36

OpenStudy (richie63):

@PMPeetaMellark Thanks! ^,^

OpenStudy (anonymous):

40/5=x x=8

OpenStudy (richie63):

5. Solve. –35 = 7t (Points : 1) –5 28 42

OpenStudy (anonymous):

-40/5=x x=-8 35=7t is equal to 35/7=t is equal to t=5

OpenStudy (anonymous):

35/7=5 5=t

OpenStudy (richie63):

Solve -5/8c =20

OpenStudy (cookiemonster2412):

divide both sides by -5/8 20/5/8=-32 c=-32

OpenStudy (richie63):

Solve -1/6b =3

OpenStudy (cookiemonster2412):

divide both sides by -1/6 b=-18

OpenStudy (richie63):

^.^ Thanks! @cookiemonster2412

OpenStudy (cookiemonster2412):

yw @Richie63